📅 Free live lesson: Tuesday 17th NovemberSave your free place →

Mixed Topic Worksheet Grades 4/5: Day 047

Grades 4/5Non-calculator
  1. Question 1. In the diagram, AB is parallel to CD.
    (a) Work out the size of the angle marked x. [1]
    (b) State the reason for your answer to part (a) [1]

    Show worked answerHide worked answer
    1. (a) B1: 112
    2. (b) B1: co-interior angles add up to 180° (the lines are parallel)
    Answer: (a) 112°; (b) Co-interior angles add up to 180°
    Co-Interior Angles
    68°xABCD
    [2 marks]
  2. Question 2. Tom flips a coin 10 times. It lands on heads 7 times.
    Tom says, "The relative frequency of heads is 0.7, so the coin is not fair."
    Is Tom correct? You must give a reason for your answer.

    Show worked answerHide worked answer
    1. B1: No
    2. B1: only 10 flips, too few to decide, so more flips are needed
    Answer: No. 10 flips is too few to tell, so he needs many more flips
    Relative Frequency Graphs
    [2 marks]
  3. Question 3. Sophie has run dfrac512 of a race.
    She runs further and has then run dfrac34 of the race.
    Work out the fraction of the race that she runs in the further part.
    Give your answer as a fraction in its simplest form.

    Show worked answerHide worked answer
    1. M1: dfrac34=dfrac912
    2. M1: 9 - 5 = 4
    3. A1: dfrac13
    Answer: dfrac13
    Subtracting Fractions
    [3 marks]
  4. Question 4. Ravi flips two fair coins.
    Ravi says, "The probability of getting two heads is dfrac12 + dfrac12 = 1."
    Is Ravi correct?
    Give a reason for your answer.

    Show worked answerHide worked answer
    1. M1: dfrac12 × dfrac12
    2. A1: No, with the probability of two heads as dfrac14
    Answer: No. The probability of two heads is 1/2 × 1/2 = 1/4, not 1.
    And/Or Rules
    [2 marks]
  5. Question 5. Here are the graphs of y = 2x - 2 and y = 10 - 2x. Jo says the solution of the simultaneous equations is x = 1 and x = 5. Is Jo correct? Give a reason.

    Show worked answerHide worked answer
    1. M1: the solution is where the lines cross, (3, 4)
    2. A1: No, x = 3 and y = 4
    Answer: No, the lines cross at x = 3, y = 4
    Solving Linear Simultaneous Equations Graphically
    xy−448y = 2x − 2y = 10 − 2x2460
    [2 marks]
📅 Free lesson: Tue 17th NovSave your place →

Mixed Topic Worksheet Grades 4/5: Day 047

Grades 4/5Non-calculator
  1. Question 1. In the diagram, AB is parallel to CD.
    (a) Work out the size of the angle marked x. [1]
    (b) State the reason for your answer to part (a) [1]

    [2 marks]
    68°xABCD
    Show worked answerHide worked answer
    1. (a) B1: 112
    2. (b) B1: co-interior angles add up to 180° (the lines are parallel)
    Answer: (a) 112°; (b) Co-interior angles add up to 180°
    Co-Interior Angles
  2. Question 2. Tom flips a coin 10 times. It lands on heads 7 times.
    Tom says, "The relative frequency of heads is 0.7, so the coin is not fair."
    Is Tom correct? You must give a reason for your answer.

    [2 marks]
    Show worked answerHide worked answer
    1. B1: No
    2. B1: only 10 flips, too few to decide, so more flips are needed
    Answer: No. 10 flips is too few to tell, so he needs many more flips
    Relative Frequency Graphs
  3. Question 3. Sophie has run dfrac512 of a race.
    She runs further and has then run dfrac34 of the race.
    Work out the fraction of the race that she runs in the further part.
    Give your answer as a fraction in its simplest form.

    [3 marks]
    Show worked answerHide worked answer
    1. M1: dfrac34=dfrac912
    2. M1: 9 - 5 = 4
    3. A1: dfrac13
    Answer: dfrac13
    Subtracting Fractions
  4. Question 4. Ravi flips two fair coins.
    Ravi says, "The probability of getting two heads is dfrac12 + dfrac12 = 1."
    Is Ravi correct?
    Give a reason for your answer.

    [2 marks]
    Show worked answerHide worked answer
    1. M1: dfrac12 × dfrac12
    2. A1: No, with the probability of two heads as dfrac14
    Answer: No. The probability of two heads is 1/2 × 1/2 = 1/4, not 1.
    And/Or Rules
  5. Question 5. Here are the graphs of y = 2x - 2 and y = 10 - 2x. Jo says the solution of the simultaneous equations is x = 1 and x = 5. Is Jo correct? Give a reason.

    [2 marks]
    xy−448y = 2x − 2y = 10 − 2x2460
    Show worked answerHide worked answer
    1. M1: the solution is where the lines cross, (3, 4)
    2. A1: No, x = 3 and y = 4
    Answer: No, the lines cross at x = 3, y = 4
    Solving Linear Simultaneous Equations Graphically
📅 Free live lesson: Tuesday 17th NovemberSave your free place →

Mixed Topic Worksheet Grades 4/5: Day 047

Grades 4/5Non-calculator
  1. Question 1. In the diagram, AB is parallel to CD.
    (a) Work out the size of the angle marked x. [1]
    (b) State the reason for your answer to part (a) [1]

    Show worked answerHide worked answer
    1. (a) B1: 112
    2. (b) B1: co-interior angles add up to 180° (the lines are parallel)
    Answer: (a) 112°; (b) Co-interior angles add up to 180°
    Co-Interior Angles
    68°xABCD
    [2 marks]
  2. Question 2. Tom flips a coin 10 times. It lands on heads 7 times.
    Tom says, "The relative frequency of heads is 0.7, so the coin is not fair."
    Is Tom correct? You must give a reason for your answer.

    Show worked answerHide worked answer
    1. B1: No
    2. B1: only 10 flips, too few to decide, so more flips are needed
    Answer: No. 10 flips is too few to tell, so he needs many more flips
    Relative Frequency Graphs
    [2 marks]
  3. Question 3. Sophie has run dfrac512 of a race.
    She runs further and has then run dfrac34 of the race.
    Work out the fraction of the race that she runs in the further part.
    Give your answer as a fraction in its simplest form.

    Show worked answerHide worked answer
    1. M1: dfrac34=dfrac912
    2. M1: 9 - 5 = 4
    3. A1: dfrac13
    Answer: dfrac13
    Subtracting Fractions
    [3 marks]
  4. Question 4. Ravi flips two fair coins.
    Ravi says, "The probability of getting two heads is dfrac12 + dfrac12 = 1."
    Is Ravi correct?
    Give a reason for your answer.

    Show worked answerHide worked answer
    1. M1: dfrac12 × dfrac12
    2. A1: No, with the probability of two heads as dfrac14
    Answer: No. The probability of two heads is 1/2 × 1/2 = 1/4, not 1.
    And/Or Rules
    [2 marks]
  5. Question 5. Here are the graphs of y = 2x - 2 and y = 10 - 2x. Jo says the solution of the simultaneous equations is x = 1 and x = 5. Is Jo correct? Give a reason.

    Show worked answerHide worked answer
    1. M1: the solution is where the lines cross, (3, 4)
    2. A1: No, x = 3 and y = 4
    Answer: No, the lines cross at x = 3, y = 4
    Solving Linear Simultaneous Equations Graphically
    xy−448y = 2x − 2y = 10 − 2x2460
    [2 marks]